Detector Characterisation (Glitches)dattilo, zaza - 17:56 Wednesday 27 August 2025 (67582)
Print this reportComment to New type of glitches around 33Hz (67574)As a follow up of the recent investigations on this new type of glitches, we made a check of the WI heating belt power supply. At around 14:57 UTC, with ITF mode set to Adjusting, we went in TCS room where the power supply is located, and also looking at its analog display, we saw that it was delivering intermittently, as indicated by the LSC_WI_HB_moni signal. We replaced the power supply with a spare one of the same kind (Kert 420). During the intervention the operator temporarily opened the WI etalon loop. It was turned back on at 15:25 UTC. Now the loop seems working regularly (attached fig). Unfortunately, during the intervention the ITF unlocked, just a few seconds after the etalon loop was turned back on.